OpenStudy (anonymous):

Some interesting integral identities. \[I_n=\int_0^\infty \exp\left(-x^n\right)\,dx=\Gamma\left(\frac{n+1}{n}\right)\\ J_n=\int_{-\infty}^\infty x^{-2n}\exp\left(-x^{-2n}\right)\,dx=\frac{1}{n}\Gamma\left(\frac{2n-1}{2n}\right)\] and one that seems related to \(J_n\), call it \({J_n}^*\), \[{J_n}^*=\int_0^\infty x^{-n}\exp\left(-x^{-n}\right)\,dx=\frac{1}{n}\Gamma\left(\frac{n-1}{n}\right)\] where \(I_n\) and \(J_n\) holds for natural \(n\ge1\), and \({J_n}^*\) for natural \(n\ge2\). To be clear, \(\large\exp(f(x))=e^{f(x)}\) and \(\Gamma\) denotes the gamma function defined by \[\Gamma(t)=\int_0^\infty x^{t-1}e^{-x}\,dx\]

OpenStudy (anonymous):

These results were obtained using Mathematica, but I'm interested in seeing an analytical approach. I attempted to verify \(I_n\) using the usual polar conversion argument for \(I_2\), which doesn't seem to work out nicely. I've also tried using the power series for \(e^t\) and integrating term by term, but I'm running into a brick wall :/ I'm currently trying out some parameterizations to see if differentiating under the integral sign might work.

OpenStudy (zarkon):

\[\Gamma\left(\frac{n+1}{n}\right)=\left(\frac{n+1}{n}-1\right)\Gamma\left(\frac{n+1}{n}-1\right)\] \[=\frac{1}{n}\Gamma\left(\frac{1}{n}\right)=\frac{1}{n}\int\limits_{0}^{\infty}u^{1/n-1}e^{-u}du\] let \(u=x^n\) \[du=nx^{n-1}dx\] \[=\frac{1}{n}\int\limits_{0}^{\infty}x^{1-n}e^{-x^n}nx^{n-1}dx\] \[=\int\limits_{0}^{\infty}e^{-x^n}dx\]

OpenStudy (zarkon):

\(J_n\) follows from starting with \(\frac{1}{n}\Gamma\left(\frac{2n-1}{2n}\right)\) use the definition of the gamma function and then do a simple substitution finally use the fact that the integrand is an even function to get the final result.

OpenStudy (anonymous):

That's quite a neat approach!

zepdrix (zepdrix):

.

OpenStudy (anonymous):

Aside from working backwards from Zarkon's method, I'm hoping to come up with a forward (left-to-right) proof of the equality. Potentially with some clever DUIS setup.

OpenStudy (kainui):

This is interesting I'll be thinking about this for a while I think to see what I can come up with.

OpenStudy (kainui):

I guess I'll just write the relationship between J and J star. \[\Large 2J_{2n}^* = J_n\] The 2n part in the subscript is a direct substitution, and the 2 multiplying the outside just comes from the fact that 2n we just plugged in makes it an even function so it's symmetrical, and half of the original J_n. I'm not sure if that's what Zarkon already discovered and explained since I'm not entirely sure I understand his argument yet. I'm playing around with the integrals more to see if there's a clever little trick.

OpenStudy (kainui):

I don't think this is what you're looking for but just starting with the definition we can work it backwards.\[\Gamma(t)=\int_0^\infty x^{t-1}e^{-x}\,dx\] \[\Gamma(\frac{n+1}{n})=\int_0^\infty x^{\frac{n+1}{n}-1}e^{-x}\,dx=\int_0^\infty x^{\frac{1}{n}}e^{-x}\,dx\]

OpenStudy (anonymous):

Right, that's what Zarkon took advantage of in his proof.

OpenStudy (kainui):

I see what you're trying to do though I think. You want to get to the answer without knowing the answer already right? Or are you more concerned with finding the method that gets this with DUIS for fun?

OpenStudy (anonymous):

Yeah I'm curious to see how one would obtain this result without knowing it beforehand. I remember seeing a tricky approach to evaluating \(I_2\) in a pdf about DUIS (probably this one: http://ocw.mit.edu/courses/mathematics/18-304-undergraduate-seminar-in-discrete-mathematics-spring-2006/projects/integratnfeynman.pdf )

OpenStudy (anonymous):

Actually, it's not in there, but I've come across another pdf very similar to it...

OpenStudy (anonymous):

Here we are, under section 4: http://www.math.uconn.edu/~kconrad/blurbs/analysis/diffunderint.pdf